# Statistical Finite Elements via Interacting Particle Langevin Dynamics

### v1

@@ Acknowledgements
Appendix A Proof of Theorem 1
@@
$\tilde{\pi}(\mathbf{b},\mathbf{p}_{1},\ldots,\mathbf{p}_{N})\propto\exp\left(-\sum_{i=1}^{N}\Phi^{y}(\mathbf{b},\sqrt{N}\mathbf{p}_{i})\right).$ (48)
It is clear that $\mathbf{b}$-marginal of this measure will concentrate on the MMAP solution as $\tilde{\pi}_{\mathbf{b}}(\mathbf{b})\propto\exp(-N\log p(\mathbf{b}|\mathbf{y}))$ by computing $\tilde{\pi}_{\mathbf{b}}(\mathbf{b})=\int\tilde{\pi}(\mathbf{b},\mathbf{p}_{1},\ldots,\mathbf{p}_{N})\mathrm{d}\mathbf{p}_{1}\ldots\mathrm{d}\mathbf{p}_{N}$. It is also clear that the potential $\Phi^{y}_{N}(\mathbf{b},\mathbf{p}_{1},\ldots,\mathbf{p}_{N})=\sum_{i=1}^{N}\Phi^{y}(\mathbf{b},\sqrt{N}\mathbf{p}_{i})$ is $N\mu$ strongly convex and $NL$ gradient Lipschitz based on Lemma 1 and [31][Lemma A.4 and A.5]. Now, let us write the standard ULA for the target (48)
$\mathrm{d}\mathbf{b}_{t}$ $=-\sum_{n=1}^{N}\nabla_{\mathbf{b}}\Phi^{y}(\mathbf{b}_{t},\sqrt{N}\mathbf{p}_{t}^{(n)})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}^{0},$
$\mathrm{d}\mathbf{p}_{t}^{(n)}$ $=-\sqrt{N}\nabla_{2}\Phi^{y}(\mathbf{b}_{t},\sqrt{N}\mathbf{p}_{t}^{(n)})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}^{n},\quad n=1,\ldots,N,$
where $\nabla_{2}$ denotes the gradient w.r.t. the second argument of $\Phi^{y}$ and $(\mathbf{B}_{t}^{n})_{t\geq 0}$ for $n=0,\ldots,N$ are $n_{u}$-dimensional Brownian motions. Consider the Euler-Maruyama discretisation of this system with a step-size $\gamma/N$
$\mathbf{b}_{k+1}$ $=\mathbf{b}_{k}-\frac{\gamma}{N}\sum_{n=1}^{N}\nabla_{\mathbf{b}}\Phi^{y}(\mathbf{b}_{k},\sqrt{N}\mathbf{p}_{k}^{(n)})+\sqrt{\frac{2\gamma}{N}}\mathbf{w}_{k+1}^{0},$
$\mathbf{p}_{k+1}^{(n)}$ $=\mathbf{p}_{k}^{(n)}-\frac{\gamma}{\sqrt{N}}\nabla_{2}\Phi^{y}(\mathbf{b}_{k},\sqrt{N}\mathbf{p}_{k}^{(n)})+\sqrt{\frac{2\gamma}{{N}}}\mathbf{w}_{k+1}^{n},\quad n=1,\ldots,N,$
where $\mathbf{w}_{k}^{n}\sim\mathcal{N}(0,I_{n_{u}})$ for every $k\geq 0$ and $n=0,\ldots,N$. It is clear that, if we write $\mathbf{u}_{k}^{(n)}=\sqrt{N}\mathbf{p}_{k}^{(n)}$, then the above system is equivalent to IPLA given in eqs. (21)–(22). We now apply Theorem 1 of [11] with $\gamma\leq 2/(\mu+L)$, which results in the bound
@@
Appendix B Forcing estimation for Linear PDE - analytic solution
To derive the MMAP forcing estimate, we start by marginalising out the latent variables $\mathbf{u}$ from the joint distribution $p(\mathbf{y}|\mathbf{b})=\int p(\mathbf{y}|\mathbf{u})p(\mathbf{u}|\mathbf{b})\mathrm{d}\mathbf{u}$. With models $\mathbf{A}\mathbf{u}=\mathbf{b}+\mathbf{e}$ and $\mathbf{y}=\mathbf{H}\mathbf{u}+\mathbf{r}$, where $\mathbf{e}\sim\mathcal{N}(\mathbf{0},\mathbf{G})$ and $\mathbf{r}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ we can marginalise analytically giving $p(\mathbf{y}|\mathbf{b})=\mathcal{N}\left(\mathbf{H}\mathbf{A}^{-1}\mathbf{b},\mathbf{H}\mathbf{A}^{-1}\mathbf{G}\mathbf{A}^{-\top}\mathbf{H}^{\top}+\mathbf{R}\right).$ With the conjugate prior $p(\mathbf{b})=\mathcal{N}(\bm{\mu},\bm{\Sigma})$, we can calculate the posterior $p(\mathbf{b}|\mathbf{y})\propto p(\mathbf{y}|\mathbf{b})p(\mathbf{b})$, where
$\log p(\mathbf{b}|\mathbf{y})=-\frac{1}{2}\left(\mathbf{y}-\mathbf{H}\mathbf{A}^{-1}\mathbf{b}\right)^{\top}\left[\mathbf{H}\mathbf{A}^{-1}\mathbf{G}\mathbf{A}^{-\top}\mathbf{H}^{\top}+\mathbf{R}\right]^{-1}\left(\mathbf{y}-\mathbf{H}\mathbf{A}^{-1}\mathbf{b}\right)$
@@ Appendix B Forcing estimation for Linear PDE - analytic solution
where MMAP estimator, $\mathbf{b}^{\star}=\operatorname*{\arg\!\max}_{\mathbf{b}}p(\mathbf{b}|\mathbf{y})$ follows as the mean of the Gaussian posterior $p(\mathbf{b}|\mathbf{y})$:
$\begin{split}\mathbf{b}^{\star}=\bigg{[}\mathbf{A}^{-\top}\mathbf{H}^{\top}\left[\mathbf{H}\mathbf{A}^{-1}\mathbf{G}\mathbf{A}^{-\top}\mathbf{H}^{\top}+\mathbf{R}\right]^{-1}\mathbf{H}\mathbf{A}^{-1}+\bm{\Sigma}^{-1}\bigg{]}^{-1}&\\ \bigg{[}\mathbf{A}^{-\top}\mathbf{H}^{\top}\left[\mathbf{H}\mathbf{A}^{-1}\mathbf{G}\mathbf{A}^{-\top}\mathbf{H}^{\top}+\mathbf{R}\right]^{-1}\mathbf{y}+&\bm{\Sigma}^{-1}\bm{\mu}\bigg{]}.\end{split}$ (50)
@@
$\mathbf{C}_{\mathbf{u}_{\mathbf{y}}}=\left[\mathbf{A}^{\top}\mathbf{G}^{-1}\mathbf{A}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right]^{-1},\quad\mathbf{m}_{\mathbf{u}|\mathbf{y}}=\mathbf{C}_{\mathbf{u}_{\mathbf{y}}}(\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{y}+\mathbf{A}^{\top}\mathbf{G}^{-1}\mathbf{b}^{\star}).$ (51)
Appendix C Nonlinear Poisson discretisation
@@
Appendix D Approximations to Nonlinearly Transformed Gaussian Random Variables
Here we present the three methods for deriving a Gaussian approximation to a nonlinearly transformed Gaussian random variable. The transform we consider is of the form $\mathbf{F}(\mathbf{u})=\mathbf{b}=\mathbf{e}$, where we wish to approximate $\mathbf{u}$ with a Gaussian, given $\mathbf{e}\sim\mathcal{N}(\mathbf{0},\mathbf{G})$.
D.1 First-order Taylor
We linearise the system about the mean point $\mathbf{m}$, found as the solution of $\mathbf{F}(\mathbf{m})=\mathbf{b}$. By replacing $\mathbf{u}_{k}$ in (58) with $\mathbf{m}$, as follows

### v2

@@ Acknowledgements
The data that support the findings of this study are available from the corresponding author, AGD, upon reasonable request.
Appendix A Proof of Lemma 1
For a symmetric block matrix $M$ of the form
$M=\begin{bmatrix}A&B\\ B^{\top}&C\end{bmatrix}$
we can use the Schur complement to determine the characteristics of this matrix in terms of its blocks, namely the fact it is positive definite (PD). We use the following from [8][Section A.5.5]:
(1) $\quad M\succ 0\iff C\succ 0\text{ and }A-BC^{-1}B^{\top}\succ 0,$
(2) $\quad\text{If }C\succ 0,\text{ then, }M\succeq 0\iff A-BC^{-1}B^{\top}\succeq 0.$
Now for the Hessian matrix we can assess if it is positive definite or positive semi definite (PSD) using these conditions, which will determine is we have $\mu>0$ for strong convexity. Define
$\nabla^{2}\Psi_{\theta}^{y}=\begin{bmatrix}\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}&-\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\\ -\mathbf{G}^{-1}\mathbf{A}_{\theta}&\mathbf{G}^{-1}+\bm{\Sigma}^{-1}\end{bmatrix}$
Now since $\mathbf{G},\bm{\Sigma}$ are covariance matrices, they are symmetric positive definite (SPD), i.e. $\mathbf{G}\succ 0,\bm{\Sigma}\succ 0$, as are their inverses $\mathbf{G}^{-1}\succ 0,\bm{\Sigma}^{-1}\succ 0$, which implies $\mathbf{G}^{-1}+\bm{\Sigma}^{-1}\succ 0$. This meets the condition of $C\succ 0$ in (1), (2). Now we can check if the Schur complement is PD or PSD which will imply PD or PSD for the Hessian $\nabla^{2}\Psi_{\theta}^{y}$. Denote by $\mathbf{S}_{\Psi}$ the Schur complement of $\nabla^{2}\Psi_{\theta}^{y}$:
$\mathbf{S}_{\Psi}$ $=\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}-\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}[\mathbf{G}^{-1}+\bm{\Sigma}^{-1}]^{-1}\mathbf{G}^{-1}\mathbf{A}_{\theta}$
$=\mathbf{A}_{\theta}^{\top}\left[\mathbf{G}^{-1}-\mathbf{G}^{-1}[\mathbf{G}^{-1}+\bm{\Sigma}^{-1}]^{-1}\mathbf{G}^{-1}\right]\mathbf{A}_{\theta}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}$
$=\mathbf{A}_{\theta}^{\top}\left[\mathbf{G}+\bm{\Sigma}\right]^{-1}\mathbf{A}_{\theta}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H},$
where we have applied Woodbury’s matrix inversion lemma
$\mathbf{G}^{-1}-\mathbf{G}^{-1}[\mathbf{G}^{-1}+\bm{\Sigma}^{-1}]^{-1}\mathbf{G}^{-1}=\left[\mathbf{G}+\bm{\Sigma}\right]^{-1},$
for the second equality. Now we show that $\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\succeq 0$ since $\mathbf{R}^{-1}$ is SPD, which means it admits a Cholesky decomposition, $\mathbf{R}^{-1}=\mathbf{L}_{\mathbf{R}}\mathbf{L}_{\mathbf{R}}^{\top}$, giving
$\mathbf{x}^{\top}\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\mathbf{x}=\mathbf{x}^{\top}\mathbf{H}^{\top}\mathbf{L}_{\mathbf{R}}\mathbf{L}_{\mathbf{R}}^{\top}\mathbf{H}\mathbf{x}=\|\mathbf{L}_{\mathbf{R}}^{\top}\mathbf{H}\mathbf{x}\|_{2}^{2}\geq 0,\quad\forall\mathbf{x}\in\mathbb{R}^{n_{u}},\mathbf{x}\neq 0.$
We can also show that $\mathbf{A}_{\theta}^{\top}[\mathbf{G}+\bm{\Sigma}]^{-1}\mathbf{A}_{\theta}$ is PD, as $\left[\mathbf{G}+\bm{\Sigma}\right]^{-1}$ is SPD and $\mathbf{A}_{\theta}$ is non-singular. We define $\mathbf{x}=\mathbf{A}_{\theta}\mathbf{y}$
$\mathbf{x}^{\top}\mathbf{A}_{\theta}^{\top}\left[\mathbf{G}+\bm{\Sigma}\right]^{-1}\mathbf{A}_{\theta}\mathbf{x}=\mathbf{y}^{\top}\mathbf{L}\mathbf{L}^{\top}\mathbf{y}=\|\mathbf{L}^{\top}\mathbf{y}\|_{2}^{2}>0,\quad\forall\mathbf{y}\in\mathbb{R}^{n_{u}},\mathbf{y}\neq 0,$
which holds because we have for every $\mathbf{x}\in\mathbb{R}^{n_{u}}$, we have $\mathbf{y}=\mathbf{A}_{\theta}^{-1}\mathbf{x}$. Putting this together, we can show the Schur complement of the Hessian is SPD:
$\mathbf{x}^{\top}\mathbf{S}_{\Psi}\mathbf{x}$ $=\mathbf{x}^{\top}\mathbf{A}_{\theta}^{\top}\left[\mathbf{G}+\bm{\Sigma}\right]^{-1}\mathbf{A}_{\theta}\mathbf{x}+\mathbf{x}^{\top}\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\mathbf{x}$
$\geq\mathbf{x}^{\top}\mathbf{A}_{\theta}^{\top}\left[\mathbf{G}+\bm{\Sigma}\right]^{-1}\mathbf{A}_{\theta}\mathbf{x}$
$>0,$
which implies $\nabla^{2}\Psi_{\theta}^{y}\succ 0$ from (1), hence $\exists\mu>0$ where $\mu=\lambda_{\text{min}}(\nabla^{2}\Psi_{\theta}^{y})$. It is also worth noting that without the inclusion of the prior on $\mathbf{b}$, when just considering the MMLE potential $\Phi_{\theta}^{y}$, we have
$\nabla^{2}\Phi_{\theta}^{y}=\begin{bmatrix}\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}&-\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\\ -\mathbf{G}^{-1}\mathbf{A}_{\theta}&\mathbf{G}^{-1}\end{bmatrix},$
and the Schur complement becomes
$\mathbf{S}_{\Phi}$ $=\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}-\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{G}\mathbf{G}^{-1}\mathbf{A}_{\theta}$
$=\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\succeq 0,$
which we have shown is only PSD, which implies via (2) that $\nabla^{2}\Phi_{\theta}^{y}\succeq 0$, and hence there may not exist a positive minimum eigenvalue of the Hessian, which only implies convexity, and not strong convexity.
Appendix B Proof of Theorem 1
@@
$\tilde{\pi}(\mathbf{b},\mathbf{p}_{1},\ldots,\mathbf{p}_{N})\propto\exp\left(-\sum_{i=1}^{N}\Psi_{\theta}^{y}(\sqrt{N}\mathbf{p}_{i},\mathbf{b})\right).$ (49)
It is clear that $\mathbf{b}$-marginal of this measure will concentrate on the MMAP solution as $\tilde{\pi}_{\mathbf{b}}(\mathbf{b})\propto\exp(N\log p(\mathbf{b}|\mathbf{y}))$ by computing $\tilde{\pi}_{\mathbf{b}}(\mathbf{b})=\int\tilde{\pi}(\mathbf{b},\mathbf{p}_{1},\ldots,\mathbf{p}_{N})\mathrm{d}\mathbf{p}_{1}\ldots\mathrm{d}\mathbf{p}_{N}$. It is also clear that the potential $\Psi^{y}_{\theta,N}(\mathbf{b},\mathbf{p}_{1},\ldots,\mathbf{p}_{N})=\sum_{i=1}^{N}\Psi^{y}_{\theta}(\sqrt{N}\mathbf{p}_{i},\mathbf{b})$ is $\tilde{\mu}=N\mu$ strongly convex and $\tilde{L}=NL$ gradient Lipschitz based on Lemma 1 and [38][Lemma A.4 and A.5]. The factor of $N$ comes as a result of the potential being a sum of $N$ potentials with Lipschitz and strong convexity constants $L$, $\mu$, and hence $N\mu\preceq\nabla^{2}\Psi_{\theta,N}^{y}\preceq NL$. Now, let us write the standard ULA for the target (49)
$\mathrm{d}\mathbf{b}_{t}$ $=-\sum_{n=1}^{N}\nabla_{\mathbf{b}}\Psi_{\theta}^{y}(\sqrt{N}\mathbf{p}_{t}^{(n)},\mathbf{b}_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}^{0},$
$\mathrm{d}\mathbf{p}_{t}^{(n)}$ $=-\sqrt{N}\nabla_{1}\Psi_{\theta}^{y}(\sqrt{N}\mathbf{p}_{t}^{(n)},\mathbf{b}_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}^{n},\quad n=1,\ldots,N,$
where $\nabla_{1}$ denotes the gradient w.r.t. the first argument of $\Psi^{y}_{\theta}$ and $(\mathbf{B}_{t}^{n})_{t\geq 0}$ for $n=0,\ldots,N$ are $n_{u}$-dimensional Brownian motions. Consider the Euler-Maruyama discretisation of this system with a step-size $\tilde{\gamma}$
$\mathbf{b}_{k+1}$ $=\mathbf{b}_{k}-\tilde{\gamma}\sum_{n=1}^{N}\nabla_{\mathbf{b}}\Psi_{\theta}^{y}(\sqrt{N}\mathbf{p}_{k}^{(n)},\mathbf{b}_{k})+\sqrt{2\tilde{\gamma}}\mathbf{w}_{k+1}^{0},$
$\mathbf{p}_{k+1}^{(n)}$ $=\mathbf{p}_{k}^{(n)}-\tilde{\gamma}\sqrt{N}\nabla_{1}\Psi_{\theta}^{y}(\sqrt{N}\mathbf{p}_{k}^{(n)},\mathbf{b}_{k})+\sqrt{2\tilde{\gamma}}\mathbf{w}_{k+1}^{n},\quad n=1,\ldots,N,$
where $\mathbf{w}_{k}^{n}\sim\mathcal{N}(0,I_{n_{u}})$ for every $k\geq 0$ and $n=0,\ldots,N$. It is clear that, if we write $\mathbf{u}_{k}^{(n)}=\sqrt{N}\mathbf{p}_{k}^{(n)}$, and set $\tilde{\gamma}=\gamma/N$ then the above system is equivalent to IPLA given in eqs. (24)–(25). We now apply Theorem 1 of [13] directly with the condition $\tilde{\gamma}\leq 2/(\tilde{\mu}+\tilde{L})$, results in the bound
$W_{2}(\nu_{k},\tilde{\pi})\leq(1-\tilde{\mu}\tilde{\gamma})^{k}W_{2}(\nu_{0},\tilde{\pi})+1.65\frac{\tilde{L}}{\tilde{\mu}}\sqrt{(N+1)n_{u}}\tilde{\gamma}^{1/2}.$
At this point, one must note the restriction can be rewritten as follows
$\tilde{\gamma}\leq 2/(\tilde{\mu}+\tilde{L})\iff\gamma\leq 2/(\mu+L).$
Hence in practice it suffices to assume $\gamma\leq 2/(\mu+L)$. Similarly, the bound above can be rewritten by substituting $\tilde{L}=NL$, $\tilde{\mu}=N\mu$, and $\tilde{\gamma}=\gamma/N$,
@@
Appendix C Forcing estimation for Linear PDE - analytic solution
To derive the MMAP forcing estimate, we start by marginalising out the latent variables $\mathbf{u}$ from the joint distribution $p(\mathbf{y}|\theta,\mathbf{b})=\int p(\mathbf{y}|\mathbf{u})p(\mathbf{u}|\theta,\mathbf{b})\mathrm{d}\mathbf{u}$. With models $\mathbf{A}_{\theta}\mathbf{u}=\mathbf{b}+\mathbf{e}$ and $\mathbf{y}=\mathbf{H}\mathbf{u}+\mathbf{r}$, where $\mathbf{e}\sim\mathcal{N}(\mathbf{0},\mathbf{G})$ and $\mathbf{r}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ we can marginalise analytically giving $p(\mathbf{y}|\theta,\mathbf{b})=\mathcal{N}\left(\mathbf{H}\mathbf{A}_{\theta}^{-1}\mathbf{b},\mathbf{H}\mathbf{A}_{\theta}^{-1}\mathbf{G}\mathbf{A}_{\theta}^{-\top}\mathbf{H}^{\top}+\mathbf{R}\right).$ With the conjugate prior $p(\mathbf{b})=\mathcal{N}(\bm{\mu},\bm{\Sigma})$, we can calculate the posterior $p(\mathbf{b}|\mathbf{y},\theta)\propto p(\mathbf{y}|\theta,\mathbf{b})p(\mathbf{b})$, where
$\log p(\mathbf{b}|\mathbf{y})=-\frac{1}{2}\left(\mathbf{y}-\mathbf{H}\mathbf{A}_{\theta}^{-1}\mathbf{b}\right)^{\top}\left[\mathbf{H}\mathbf{A}_{\theta}^{-1}\mathbf{G}\mathbf{A}_{\theta}^{-\top}\mathbf{H}^{\top}+\mathbf{R}\right]^{-1}\left(\mathbf{y}-\mathbf{H}\mathbf{A}_{\theta}^{-1}\mathbf{b}\right)$
@@ Appendix C Forcing estimation for Linear PDE - analytic solution
where MMAP estimator, $\mathbf{b}^{\star}=\operatorname*{\arg\!\max}_{\mathbf{b}}p(\mathbf{b}|\theta,\mathbf{y})$ follows as the mean of the Gaussian posterior $p(\mathbf{b}|\theta,\mathbf{y})$:
$\begin{split}\mathbf{b}^{\star}=\bigg{[}\mathbf{A}_{\theta}^{-\top}\mathbf{H}^{\top}\left[\mathbf{H}\mathbf{A}_{\theta}^{-1}\mathbf{G}\mathbf{A}_{\theta}^{-\top}\mathbf{H}^{\top}+\mathbf{R}\right]^{-1}\mathbf{H}\mathbf{A}_{\theta}^{-1}+\bm{\Sigma}^{-1}\bigg{]}^{-1}&\\ \bigg{[}\mathbf{A}_{\theta}^{-\top}\mathbf{H}^{\top}\left[\mathbf{H}\mathbf{A}_{\theta}^{-1}\mathbf{G}\mathbf{A}_{\theta}^{-\top}\mathbf{H}^{\top}+\mathbf{R}\right]^{-1}\mathbf{y}+&\bm{\Sigma}^{-1}\bm{\mu}\bigg{]}.\end{split}$ (51)
@@
$\mathbf{C}_{\mathbf{u}|\mathbf{y}}=\left[\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right]^{-1},\quad\mathbf{m}_{\mathbf{u}|\mathbf{y}}=\mathbf{C}_{\mathbf{u}|\mathbf{y}}(\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{y}+\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\mathbf{b}^{\star}).$ (52)
Appendix D Nonlinear Poisson discretisation
@@
Appendix E Approximations to Nonlinearly Transformed Gaussian Random Variables
Here we present the three methods for deriving a Gaussian approximation to a nonlinearly transformed Gaussian random variable. The transform we consider is of the form $\mathbf{F}(\mathbf{u})=\mathbf{b}+\mathbf{e}$, where we wish to approximate $\mathbf{u}$ with a Gaussian, given $\mathbf{e}\sim\mathcal{N}(\mathbf{0},\mathbf{G})$.
E.1 First-order Taylor
We linearise the system about the mean point $\mathbf{m}$, found as the solution of $\mathbf{F}(\mathbf{m})=\mathbf{b}$. By replacing $\mathbf{u}$ in (59) with $\mathbf{m}$, as follows